# Augmenting Replay in World Models for Continual Reinforcement Learning

### 30 Jan 2024 v1

@@ 3.1.1 Baseline Normalisation
We perform 2 evaluations of each environment to be considered as ‘baseline’ performance. For each task in the experimental suite $\tau\in\boldsymbol{\mathcal{T}}$, we find the average episodic reward of a random policy $p^{(\tau)}_{\tau}(0)$ and the average episodic reward of the agent on only task $\tau$ after $N$ environment steps $p^{(\tau)}_{\tau}(N)$. The normalized score of subsequent agents will be such that a score of 00 is equal to random performance, and a score of $1$ is equal to the final performance.
@@
Average random performance and average final performance of WMAR, when trained on a single task, are used for score normalisation. A score equal to the random performance would correspond to 0, and a score corresponding to the average final performance would correspond to 1.
@@ 3.1.2 Continual Learning Assessment
We follow Kessler et al. (2022) in evaluating the efficacy of a continual learning method in forgetting, measuring stability and backward transfer, and forward transfer, also acting as a measure for plasticity. These measurements are recorded for each environment in the experiment suite $\boldsymbol{\mathcal{T}}=(\tau_{1},\tau_{2},\dots,\tau_{T})$. We train the agent on each task for $N$ environment steps. As such, the reference score for task $\tau_{i}$ would be given by $q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(i\times N)$.
Forgetting
@@
A lower value for forgetting is indicative of improved stability and a superior continual learning method. A negative value for forgetting would imply that the agent has managed to gain performance on earlier tasks, thus exhibiting backward transfer.
@@
The larger the forward transfer, the better the continual learning method. A positive value for forward transfer would imply that the system effectively uses learnt knowledge from previous environments and has accelerated learning in the current environment as a result. When each task is not related to each other, no positive forward transfer may be expected. In this case, a forward transfer of 0 would represent optimal plasticity, and negative values would indicate a barrier to learning newer tasks by previous tasks.
@@ 3.2 Environments
We explore the performance of WMAR within the context of transfer between tasks with and without shared structure, which we define as two tasks having similar environment state and observations, action dynamics, and rewards. To investigate the performance on tasks with shared structure, we use OpenAI Procgen’s CoinRun environment and apply visual perturbations to distinguish between different environments. For evaluating the performance on tasks without shared structure, we use 4 different tasks from Atari.
@@ 4.1 Tasks with Shared Structure
The CoinRun environment provides a number of simple visual perturbations. While the environment’s mechanics remain identical, the environments share both subtle and significant visual differences, which can pose significant risks to CL characteristics. However, we show that WMAR enables constructive interference and positive forward transfer of skills between these environments.
Baseline Performance
We measure the performance of WMAR and other baseline RL algorithms when applied to each task within the suite individually (see Section S.LABEL:S-sec:baseline for parameters and results). The performance of WMAR is approximately 15% better across tasks compared to the PPO (CNN-LSTM) algorithm, but worse compared to DreamerV3, likely due to implementation differences. The ‘Random’ and ‘WMAR’ scores are used as the normalisation baselines for each task.
@@
The normalised performance is shown in Figure 3 (tabular results in Section LABEL:S-sec:cl). The terminology that denotes the variant of the CoinRun environment is as follows. NB = no background, RT = restricted themes, and MA = Monochrome assets (see Section S.LABEL:S-sec:procgen).
@@
Regardless of the choice of replay buffer, the agent is capable of exhibiting desirable forgetting and backward transfer properties. However, we observe a slight benefit to those desiderata when the long-term global distribution matching buffer is used in conjunction with the FIFO buffer, evidenced by the improved forgetting and backward transfer (Table 1). For the FIFO only implementation, no backward transfer on the ‘CoinRun’ environment is observed. This contrasts to the backward transfer observed for the same environment where the distribution matching buffer is present (Section S.LABEL:S-tab:raw_res_coinrun). Further, we observe that the method using the distribution matching buffer shows more consistent performance across tasks with fewer fluctuations in performance during training (Figure 3), which may be a desirable property if there exists a need to snapshot and use the agent at arbitrary points during training.
@@
Our results show no evidence of inhibition to the strength of the agent when it is required to learn multiple tasks with shared structure. Rather, we find little difference on the resulting forward transfer regardless of the inclusion of the distribution matching buffer, see Table 1. We find significant forward transfer on all tasks after the first, greatly lowering the amount of time required for a task to reach a certain performance. Importantly, the ordering of the tasks is such that levels in later tasks are either subsets of, or different to levels in earlier tasks. This is so that the agent is unable to completely generalise on earlier tasks just by learning later tasks. Crucially, our results show that strong forward transfer also holds for tasks that are not subsets of earlier tasks.
Figure 3: Continual learning performance of WMAR with (top) or without (bottom) long-term distribution matching buffer, with bold line segments denoting the periods in which certain tasks are being trained. Scores are normalised by Equation 1. The line is the median and the shaded area is between the 25th and 85th percentile, of 5 seeds.
@@
4.2 Tasks without Shared Structure
We use a subset of Atari environments in which the world model agent can achieve reasonable performance with little training and follow Machado et al. (2018) in using sticky actions in Atari environments. We run the CL experiment on an arbitrary ordering of the environments as testing each permutation would be prohibitively expensive. We show that significant divergence between environment dynamics in visual differences, reward scale, and game mechanics pose a significant challenge to CL. Moreover, lacking a task oracle, we find predetermined reward scales (Table S.LABEL:S-tab:atari_rew_scales) for each individual environment necessary to enable continual learning.
Baseline Performance
WMAR’s performance remains competitive against other common RL algorithms on single tasks (see Section S.LABEL:S-sec:baseline for parameters and results). Again, the ‘Random’ and ‘WMAR’ scores are used as the normalisation baselines for each task.
Continual Learning
The normalised performance is plotted in Figure 4 (tabular results in Section LABEL:S-sec:cl). It is clear that DreamerV3 is not capable of CL; it only performs well when it is actively training on a given task.
Forgetting
WMAR shows significantly improved stability in learning at all points after the task is learned. With only the FIFO buffer, and in all tasks except the last, the agent loses virtually all performance before the following task has finished training. This is in direct contrast to WMAR with the addition of the long-term global distribution matching buffer where instead, the agent is able to consistently maintain much of its performance on tasks learned earlier on (Table 2). This difference in forgetting emphasises that the distribution matching buffer is necessary to prevent forgetting when tasks are significantly different, as the learning of new tasks causes knowledge of previously learnt tasks to be lost from the world model. As training progresses, samples from the first task have a decreasing probability of being drawn. This is to the point where, by the end of training, samples from the first task have a 12.5% probability of being drawn. Despite this, our results show that WMAR is able to hold much of its performance on the earliest of tasks (Figure 4). Notably, this observation does not apply to all tasks, as we find that the ‘Crazy Climber’ task exhibits the worst overall performance and forgetting even with the distribution matching buffer.
Figure 4: Continual learning performance of WMAR with (top) or without (middle) long-term distribution matching buffer, with bold line segments denoting the periods in which certain tasks are being trained. DreamerV3 benchmark results are also given (bottom). Scores are normalised by Equation 1. The line is the median and the shaded area is between the 25th and 85th percentile, of 5 seeds.
Model Avg. forgetting Avg. fwd. transfer
(Lower is better)
(Higher is better)
WMAR
$\mathbf{0.10}\pm 0.04$
$-0.35\pm 0.10$
WMAR (FIFO)
$0.68\pm 0.10$
$\mathbf{-0.11}\pm 0.14$
Forward Transfer and Plasticity
We find significant differences in the properties of continual learning on tasks with and without shared structure which are reflected in the results. Contrary to findings in tasks with shared structure, the continual learning setting offers no benefit to forward transfer. Benefits to continual learning in greatly reduced forgetting come at the cost of also reduced forward transfer — effectively trading some plasticity for a significant amount of stability, as the agent is unable to learn later tasks as effectively. We observe a decreasing trend in the agent’s normalized score as it learns more tasks, most apparent in the WMAR experiment, where the agent achieves a normalised score of only $0.49\pm 0.09$ in the ‘Crazy Climber’ environment, substantially lower than the expected 1.0 of the first environment ‘Ms Pacman’ and empirical $0.84\pm 0.04$ of ‘Boxing’ (Table S.LABEL:S-tab:raw_res_atari), which is already well below 1.0. This evidences that it performs worse when learning new tasks after learning unrelated tasks, suggesting that the system is limited by its learning capacity in either the world model or the agent. This limitation thus results in plasticity being gradually lost as training progresses. Moreover, the negative forward transfer shown by experiments with and without the distribution matching buffer suggests that the agent learns newer tasks slower when it has already learned a previous task as compared to if it were randomly initialised, further suggesting that plasticity is a practical constraint in CL with world models.
@@ 5 Discussion
WMAR has the ability to perform CL with reduced memory requirements. On tasks with shared structure, it exhibits improved forward and backward transfer, and on tasks without shared structure, it exhibits reduced forgetting. However, fully achieving all desiderata of stability and plasticity under all conditions (with or without shared structure), remains a challenge. Specifically, we show that tasks without shared structure, lacking visual similarity and having different mechanics, may be learned less effectively if they are learned later.
@@
Overall, these results were made possible by taking advantage of a distribution matching buffer running alongside a typical FIFO buffer. This change, coupled with the world model approach to RL, enables improved sample efficiency and can operate with significantly smaller replay buffers. Further, being able to learn tasks quickly benefits scalability of the algorithm.
@@
Our results reiterate findings of world model approaches as strong baselines for CL problems, which we attribute to their natural ability to support off-policy experience replay. Our methods are also largely orthogonal to previous state-of-the-art approaches to combating catastrophic forgetting such as EWC and P&C which work over network parameters, and CLEAR which utilises replay but typically operates over model-free approaches and uses behaviour cloning of the policy from environment input to action output. Yet, we show that WMAR remains capable of eliminating catastrophic forgetting across a diverse set of environments.
5.1 Generalisation of Tasks
@@
While we find tasks with shared structure to enable positive forward and backward transfer, we note that this is not the case in tasks without shared structure. We posit this is due to the reliance on the convolutional feature extractor within the WMAR architecture, enabling it to only take advantage of commonality in the visual domain. Much more abstract similarities or changes in the controls between tasks are likely to negatively impact the forward transfer, causing lower average and maximum performance of later tasks. For example, a change to an environment involving inverting the colours and permuting the controls is likely to induce negative forward transfer. A key distinction between CL with world models and model-free methods is the requirement to reconstruct the high-dimensional observations in the form of images. We posit this induces significant strain on the capacity of the model, possibly in the learning of the latent space, or in the model parameters for encoding and reconstructing the observation, especially if it varies greatly between tasks.
5.2 Reward Scaling without Task Oracle
@@
We also find that approximate reward scaling is required to learn multiple tasks which naturally have extremely different reward scales. While WMAR is sufficiently robust in learning multiple tasks without forgetting even when reward scales are somewhat different, we find this property to not hold when reward scales differ significantly (e.g. by a factor of $10^{2}$), where only the task with the higher reward scale will be learnt. Barring the possibility of access to a task oracle, we believe that this limitation affects all continual learning applications. This limitation is likely a result of the different rewards and subsequent returns causing poorly scaled advantages when training the actor, resulting in the actor only learning tasks with the highest return magnitudes. Experiments involving approaches with automatic scaling of advantages through non-linear squashing transformations proved to hurt learning on individual tasks, so a static, linear, reward transformation was used.
5.3 Limitations of Reservoir Sampling
@@
The WMAR approach brings substantial improvement to memory requirements and CL characteristics, all while significantly reducing catastrophic forgetting. However, the scalability of this approach to a larger number of tasks remains unproven. A key limitation inherited from the use of reservoir sampling is the finite capacity of the long-term global distribution matching buffer. As more tasks are explored and previous tasks are not revisited, an increasing number of samples from previous tasks will inevitably be lost, leading to forgetting. Indeed, we observe some variance in the amount of forgetting to which each task is susceptible. Results from the experiments on tasks without shared structure (Figure 4) suggest latter tasks which may have been learned poorly due to insufficient plasticity in the learning system may experience more forgetting than even earlier tasks.
A plausible approach to improving the performance of WMAR on a larger number of tasks is to increase the size of the long-term global distribution matching buffer. However, this approach naturally inherits limitations from world model approaches and memory-based CL approaches, as doing so incurs significant memory requirements. While techniques such as image compression may alleviate this constraint, RL is frequently done at scale, where the need to send and store a significant volume of image data may pose a significant engineering challenge, reducing the applicability and accessibility of this algorithm.

### 05 Mar 2024 v2

@@ 3.1.1 Baseline Normalisation
We performed 2 evaluations of each environment to be considered as ‘baseline’ performance. For each task in the experimental suite $\tau\in\boldsymbol{\mathcal{T}}$, we find the average episodic reward of a random policy $p^{(\tau)}_{\tau}(0)$ and the average episodic reward of the agent on only task $\tau$ after $N$ environment steps $p^{(\tau)}_{\tau}(N)$. The normalised score of subsequent agents will be such that a score of 00 is equal to random performance, and a score of $1$ is equal to the final performance.
@@
Average random performance and average final performance of WMAR, when trained on a single task, were used for score normalisation. A score equal to the random performance corresponds to 0, and a score corresponding to the average final performance corresponds to 1.
@@ 3.1.2 Continual Learning Assessment
We followed Kessler et al. (2022) in evaluating the efficacy of a continual learning method in forgetting, measuring stability and backward transfer, and forward transfer, also acting as a measure for plasticity. These measurements were recorded for each environment in the experiment suite $\boldsymbol{\mathcal{T}}=(\tau_{1},\tau_{2},\dots,\tau_{T})$. We trained the agent on each task for $N$ environment steps. As such, the reference score for task $\tau_{i}$ is given by $q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(i\times N)$.
Forgetting and Backward Transfer
@@
A lower value for forgetting is indicative of improved stability and a better continual learning method. A negative value for forgetting would imply that the agent has managed to gain performance on earlier tasks, thus exhibiting backward transfer.
@@
The larger the forward transfer, the better the continual learning method. A positive value for forward transfer implies that the system effectively uses learnt knowledge from previous environments and has accelerated learning in the current environment as a result. When each task is not related to the others, no positive forward transfer may be expected. In this case, a forward transfer of 0 would represent optimal plasticity, and negative values would indicate a barrier to learning newer tasks from previous tasks.
@@ 3.2 Environments
We explored the performance of WMAR within the context of transfer between tasks with and without shared structure, which we define as two tasks having similar environment state and observations, action dynamics, and rewards. To investigate the performance on tasks with shared structure, we use OpenAI Procgen’s CoinRun environment and apply visual perturbations to create differences between tasks. For evaluating the performance on tasks without shared structure, we selected 4 Atari tasks.
@@ 4 Results
4.1 Tasks without Shared Structure
We chose a subset of Atari environments in which the world model agent can achieve reasonable performance with little training and followed Machado et al. (2018) in using sticky actions. We ran the CL experiment with arbitrary ordering of the environments as testing each permutation would be prohibitively expensive. We found that significant divergence between environment dynamics in visual differences, reward scale, and game mechanics pose a significant challenge to CL. Furthermore, without a task oracle, we found predetermined reward scales (Table 8) for each individual environment necessary to allow continual learning.
Baseline Performance (Single Tasks)
WMAR’s performance remains competitive against other common RL algorithms on single tasks (see Appendix D.2 for parameters and results). The ‘Random’ and ‘WMAR’ scores are used as normalisation baselines for each task.
Continual Learning
The normalised performance is plotted in Figure 3 (tabular results in Section D.3, Table 8). WMAR is capable of continual RL, whereas it is clear that WMAR (FIFO only) the DreamerV3 equivalent, is not; it only performs well when it is actively training on a given task.
Forgetting and Backward Transfer
In all tasks except the last, WMAR (FIFO only) loses virtually all performance before the following task has finished training. This is in direct contrast to WMAR, where instead the agent can maintain much of its performance on previously learnt tasks, see Table 1. In summary, WMAR shows significantly improved stability. This difference in forgetting emphasises that the distribution matching buffer is necessary to prevent forgetting when tasks are significantly different, as the learning of new tasks causes knowledge of previously learnt tasks to be lost from the world model. As training progresses, samples from the first task have a decreasing probability of being drawn. This is to the point where, by the end of training, samples from the first task have a 12.5% probability of being drawn. Despite this, our results show that WMAR is capable of holding much of its performance on the earliest of tasks. Notably, this observation does not apply to all tasks, as we find that the ‘Crazy Climber’ task exhibits the worst overall performance and forgetting even with the distribution matching buffer.
Forward Transfer
Reduced forgetting comes at the cost of also reduced forward transfer – effectively trading some plasticity for a significant amount of stability – as the agent is unable to learn later tasks as effectively. We observe a decreasing trend in the agent’s normalised score as it learns more tasks, most apparent in the WMAR experiment, where the agent achieves a normalised score of only $0.49\pm 0.09$ in the ‘Crazy Climber’ environment, substantially lower than the expected 1.0 of the first environment ‘Ms Pacman’ and empirical $0.84\pm 0.04$ of ‘Boxing’, which is already well below 1.0. The system performs worse when learning new tasks after learning unrelated tasks, suggesting that the system is limited by its learning capacity in either the world model or the policy. This limitation thus results in plasticity being gradually lost as training progresses. Furthermore, the negative forward transfer shown by experiments with and without the distribution matching buffer suggests that the agent learns newer tasks slower when it has already learnt a previous task compared to if it were randomly initialised, further suggesting that plasticity is a practical constraint in continual RL with world models.
Figure 3: Continual learning performance on tasks without shared structure. Bold line segments denote periods in which certain tasks are trained. Scores are normalised by Equation 1. The line is the median and the shaded area is between the 25th and 85th percentile, of 5 seeds.
Model Avg. forgetting Avg. fwd. transfer
(Lower is better)
(Higher is better)
WMAR
$\mathbf{0.10}\pm 0.04$
$-0.35\pm 0.10$
WMAR (FIFO)
$0.68\pm 0.10$
$\mathbf{-0.11}\pm 0.14$
@@ 4.2 Tasks with Shared Structure
The CoinRun environment provides a number of simple visual perturbations. While the environment’s mechanics remain identical, the environments share both subtle and substantial visual differences, which poses significant challenges. However, we found that WMAR enables constructive interference and positive forward transfer of skills between these environments.
Baseline Performance – Single Tasks
The performance of WMAR is approximately 15% better in all tasks compared to the PPO (CNN-LSTM) algorithm, but worse compared to DreamerV3, likely due to implementation differences. Again, the ‘Random’ and ‘WMAR’ scores are used as the normalisation baselines for each task. See Appendix D.2 for parameters and detailed results.
@@
The normalised performance is shown in Figure 4 (tabular results in Appendix D.3, Table 10). The terminology that denotes the variant of the CoinRun environment is as follows. NB = no background, RT = restricted themes, and MA = monochrome assets (see Appendix C.1). Both models can learn continuously and improve in performance and have approximately equal peak performance on each task. However, WMAR is noticeably more consistent than WMAR (FIFO Only).
@@
Regardless of the choice of replay buffer, the agent exhibits desirable forgetting and backward transfer. However, WMAR has a slight advantage over WMAR (FIFO only), where almost no backward transfer is observed; see Table 2. Further, WMAR shows more consistent performance across tasks with fewer fluctuations in performance during training, which may be a desirable property if there is a need to take a model snapshot and use the agent at arbitrary points during training.
@@
There is no evidence of inhibition to the strength of the agent when learning multiple tasks. Rather, we find little difference on the resulting forward transfer regardless of the inclusion of the distribution matching buffer; see Table 2. We found significant forward transfer on all tasks after the first, greatly lowering the amount of time required for a task to reach a certain performance. Importantly, the ordering of the tasks is such that levels in later tasks are either subsets of, or different to levels in earlier tasks. This is so that the agent cannot completely generalise on earlier tasks just by learning later tasks. Crucially, our results show that strong forward transfer also holds for tasks that are not subsets of earlier tasks.
Figure 4: Continual learning performance on tasks with shared structure. Bold line segments denote the periods in which certain tasks are trained. Scores are normalised by Equation 1. The line is the median and the shaded area is between the 25th and 85th percentile, of 5 seeds.
@@ 5 Discussion
WMAR has the ability to perform CL with reduced memory requirements compared to DreamerV3. On tasks with shared structure, it exhibits improved forward and backward transfer, and on tasks without shared structure, it exhibits dramatically reduced forgetting. However, fully achieving all desiderata of stability and plasticity under all conditions (with or without shared structure) remains a challenge. Specifically, we show that tasks without shared structure, lacking visual similarity and having different mechanics, may be learnt less effectively if they are learnt later.
@@
Overall, these results were made possible by taking advantage of a distribution matching buffer running alongside a typical FIFO buffer. This change, coupled with the world model approach to RL, enables improved sample efficiency and can operate with significantly smaller replay buffers.
@@
Our results reiterate findings of world model approaches as strong baselines for continual RL problems, which we attribute to their natural ability to support off-policy experience replay. Our methods are also largely orthogonal to previous state-of-the-art approaches to combating catastrophic forgetting such as EWC and P&C which work over network parameters, and CLEAR which utilises replay but typically operates over model-free approaches and uses behaviour cloning of the policy from environment input to action output. Yet, we show that WMAR remains capable of eliminating catastrophic forgetting across a diverse set of environments.
Generalisation of Tasks
@@
While we observed positive forward and backward transfer on tasks with shared structure, it was not the case for tasks without. We posit this is due to reliance on the convolutional feature extractor, which could benefit from commonality in the visual domain only. Much more abstract similarities or changes in the controls between tasks are likely to negatively impact the forward transfer, causing lower average and maximum performance of later tasks. For example, changes such as inverting the colours and permuting the controls is likely to induce negative forward transfer.
Reward Scaling without Task Oracle
@@
We found that approximate reward scaling is required to learn multiple tasks which naturally have extremely different reward scales. While WMAR is sufficiently robust in learning multiple tasks without forgetting even when reward scales are somewhat different, we found that this property did not hold when reward scales differed significantly (e.g. by a factor of $10^{2}$), where only the task with the higher reward scale will be learnt. We believe that without a task oracle, this limitation affects all continual RL. It is likely a result of the different rewards and subsequent returns causing poorly scaled advantages when training the actor, resulting in the actor only learning tasks with the highest return magnitudes. Experiments with automatic scaling of advantages through non-linear squashing transformations proved to hurt learning on individual tasks, so a static, linear, reward transformation was used.
Memory Capacity
@@
Despite the benefits and improved memory capacity of WMAR, a key limitation of any buffer-based method is finite capacity. In our case, as more tasks are explored and previous tasks are not revisited, an increasing number of samples from previous tasks will inevitably be lost, leading to forgetting. Indeed, we observe some variance in the amount of forgetting to which each task is susceptible. Results from the experiments on tasks without shared structure (Figure 3) suggest that the latter tasks which may have been learnt poorly due to insufficient plasticity in the learning system may experience more forgetting than even earlier tasks.
@@ 6 Conclusion and Future Work
We extended a well-known model-based ‘world model’ architecture, DreamerV3, with an improved memory buffer, and applied it to the problem of continual RL in two settings: a continuous stream of tasks with and without shared structure i.e. commonalities between the tasks that could be leveraged by the agents. Performance was evaluated for forward and backward transfer, in addition to the common practice of measuring only forgetting. We found that model-based agents are capable of continual learning on tasks with shared structure, with some minor benefits to the enhanced buffer in WMAR. And in tasks without shared structure, the improved buffer in WMAR conferred substantial improvements. In both cases, WMAR has substantially reduced memory requirements. The results suggest that model-based RL using a world model with a memory-efficient replay buffer can be an effective and practical approach to continual RL, justifying future work.